# The Linear Slicing Method for Equal Sums of Like Powers: Modular and Geometric Constraints

### v1

@@ Abstract
with integer variables and real exponent $k>1$, under the additional linear constraint
@@
On the central slice $h=0$ we show, using strict convexity, that there are no non-trivial solutions even over the reals: if $a,b,c,d\geq 0$ and $a+b=c+d$, then $a^{k}+b^{k}=c^{k}+d^{k}$ forces $\{a,b\}=\{c,d\}$. For integer $k\geq 2$ we also prove a quantitative separation theorem on the hyperplane $a+b=c+d$, showing that distinct unordered pairs with the same sum $S$ produce values of $a^{k}+b^{k}$ that are separated by $\gg_{k}S^{k-2}$.
@@
Then any integer solution with shift $h$ must satisfy $M_{k}\mid h$; this modulus $M_{k}$ is the maximal squarefree modulus for which $x^{k}\equiv x\pmod{M_{k}}$ holds for all integers $x$. As immediate corollaries, if $k$ is prime then $k\mid h$, and for fixed $k$ the necessary condition $M_{k}\mid h$ leaves only a $1/M_{k}$ fraction of shifts. In particular, this creates a divisibility filter that eliminates, for instance, $99.96\%$ of all linear slices for $k=13$ (since $M_{13}=2730$). Combining MDO with our convexity analysis yields an exclusion zone principle. While this implies a synergy bound on slice size $S\geq 2h/(k-1)$, we show that a global overlap argument provides a strictly stronger constraint. Finally, along any fixed slice $(S,h)$ we prove an asymptotic dominance bound $k\leq\max\{S,S+h\}\log 2$, beyond which no integer solutions can occur.
@@ 1 Introduction
For $k\geq 5$, however, the situation is dramatically different. It is a well-known open problem whether there exists any non-trivial integer solution to (1) in this range; see, for example, Browning [2] and the references therein, as well as Guy’s collection of unsolved problems [3].
@@
  • for integer $k\geq 2$, an exclusion zone of radius $\asymp\sqrt{S}$ around $S/2$ on the smaller-sum slice $a+b=S$ for each fixed $h\neq 0$;
@@
Combining MDO with the exclusion zone gives the lower bound $S\geq 2M_{k}/(k-1)$ whenever $h\neq 0$.
Organization of the paper. Section 2 develops the theory of the central slice $h=0$, including qualitative uniqueness and quantitative separation on $a+b=c+d$. Section 3 formulates central-slice uniqueness for general strictly convex functions. Section 4 proves the modular divisibility obstruction (MDO), its maximality, and corollaries (parity as a special case, density, prime exponents), and includes a table of $M_{k}$ for odd $k$. Section 5 contains the exclusion zone principle and its synergy with MDO, yielding $S\geq 2M_{k}/(k-1)$ for $h\neq 0$, together with a separate “overlap” bound coming from the global ranges of $a^{k}+b^{k}$ on different slices. Section 6 establishes the asymptotic dominance bounds for fixed slices. Finally, Section 7 discusses heuristics and broader context.
Throughout, $\mathbb{N}=\{0,1,2,\dots\}$ and we mostly work with non-negative variables $a,b,c,d\geq 0$; this entails no loss of generality for our purposes. Unless explicitly stated, $\log$ denotes the natural logarithm.
@@ 2.2 Quantitative separation on the central slice
Apply Lemma 2.5 to $f_{S}$ and bound the last difference on $0,\lfloor S/2\rfloor]$ from below as in the standard binomial expansion argument (even/odd $S$ cases). ∎
@@ 4 Modular constraints on the shift: the MDO
If $p\mid N$, then $x^{k}\equiv x\pmod{p}$ for all $x$. Restricting to units shows $x^{k-1}\equiv 1$ for all $x\in(\mathbb{Z}/p\mathbb{Z})^{\times}$, so the exponent of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ divides k1k-1. But that group has order p1p-1; hence $p-1\mid k-1$, i.e. $p\in\mathcal{P}_{k}$, and therefore $p\mid M_{k}$. As $N$ is squarefree, $N\mid M_{k}$. ∎
@@
For odd $k\in\{3,5,\dots,19\}$ we record $\mathcal{P}_{k}$, the squarefree modulus $M_{k}$, the density $1/M_{k}$, and the synergy lower bound on the slice size (see Proposition 5.5) written as $2M_{k}/(k-1)$.
@@ 19 $\{2,3,7,19\}$ $798$ $1/798$ $1596/18$
While the table lists $M_{k}$ for $k\leq 19$, the value of $M_{k}$ grows explosively for exponents with highly composite k1k-1. For instance, at $k=61$ we have $60=2^{2}\cdot 3\cdot 5$, leading to
@@ 5 Geometry of shifted slices and exclusion zones
We will prove the following inequality, which shows that the $k$-th power sum is bounded above by the $k$-th power of its quadratic approximation at the mean $X$:
@@
Since the function $t\mapsto t^{k}$ is strictly increasing for $t\geq 0$, taking the $k$-th root yields:
@@
(For $m=0$, both sides are 1. For $m>\lfloor k/2\rfloor$, the LHS coefficient is 0, while RHS is non-negative.)
Case $m=1$: LHS is $\binom{k}{2}=\frac{k(k-1)}{2}$. RHS is $\binom{k}{1}\frac{k-1}{2}=\frac{k(k-1)}{2}$. They are equal.
Case $m\geq 2$: We rewrite (4) (assuming $2m\leq k$):
@@
We show that each factor in this product is less than or equal to (k1)/2(k-1)/2. We check the inequality:
$\frac{k-m-j}{2m-j}\leq\frac{k-1}{2}$
$2(k-m-j)\leq(k-1)(2m-j)$
$2k-2m-2j\leq 2mk-2m-jk+j$
$2k(1-m)\leq j(3-k).$
Since $m\geq 2$, the LHS is negative. If $k=2$ or $k=3$, the condition $2m\leq k$ implies $m\leq 1$, contradicting $m\geq 2$. So this case is empty for $k=2,3$. If $k\geq 4$. We require $2k(m-1)\geq j(k-3)$ (multiplying by 1-1). Since $j\leq m-1$ and k3>0k-3>0, it suffices to show $2k(m-1)\geq(m-1)(k-3)$. Since $m>1$, we need $2k\geq k-3$, or $k\geq-3$, which is true.
Thus, every factor in the LHS product is less than or equal to (k1)/2(k-1)/2, so the entire product is less than or equal to $((k-1)/2)^{m}$.
@@
Proposition 5.5 (Synergy with MDO: lower bound on slice size).
@@
From $\delta\leq S/2$ and Theorem 5.4 we get $S^{2}/4\geq\delta^{2}\geq Sh/(2(k-1))$, which implies $S\geq 2h/(k-1)$. The second bound follows since $|h|\geq M_{k}$ when $h\neq 0$ and $M_{k}\mid h$. ∎
@@
It is important to note that the overlap bound (Lemma 5.2) is strictly stronger than the synergy bound (Proposition 5.5) for all $k\geq 2$. That is,
@@ 6 Asymptotic dominance on fixed slices
If the maximal base $M$ appeared on both sides, then $h=0$. Hence for $h\neq 0$ it appears on exactly one side; say $M^{k}+b^{k}=c^{k}+d^{k}$ with $b,c,d\leq M-1$. Then $M^{k}-(M-1)^{k}\leq|d^{k}-b^{k}|\leq(M-1)^{k}$, which fails when $k>M\log 2$ by Lemma 6.1. Thus $k\leq M\log 2\leq S_{0}\log 2$. ∎
@@ 7 Concluding remarks
On the central slice $a+b=c+d$ we have complete uniqueness (Theorem 2.1) and quantitative separation (Theorem 2.6): distinct unordered pairs $\{a,b\}$ along $a+b=S$ produce values of $a^{k}+b^{k}$ separated by $\gg_{k}S^{k-2}$. This gives a simple local avatar of sparsity heuristics for equal sums of two $k$th powers.
For shifted slices $h\neq 0$, MDO (Theorem 4.1) produces a squarefree modulus $M_{k}$ with $M_{k}\mid h$, maximal among such moduli by Lemma 4.2. This contains parity as a special case (Corollary 4.3), yields $k\mid h$ for prime $k$, and gives density $1/M_{k}$ of admissible shifts (Corollary 4.5). For many odd $k$ (e.g. $k=13$) this drastically thins the set of feasible $h$. Combined with the exclusion zone (Theorem 5.4) we obtain the lower bound $S\geq 2M_{k}/(k-1)$ (Proposition 5.5), which is a clean structural constraint on the slice size. In addition, the overlap bound of Lemma 5.2 shows that the smaller of the two slice sums must satisfy $\min\{S,S+h\}\gg_{k}|h|$, giving a global geometric restriction based solely on the ranges of $a^{k}+b^{k}$ on the two slices. Finally, along a fixed slice $(S,h)$, the dominance bound $k\leq\max\{S,S+h\}\log 2$ (Theorem 6.2) shows that only finitely many exponents can occur.

### v2

@@ Abstract
with non-negative integer variables $a,b,c,d\in\mathbb{Z}_{\geq 0}$ and real exponent $k>1$, under the additional linear constraint
@@
On the central slice $h=0$, we show, using strict convexity, that there are no non-trivial solutions even over the reals: if $a,b,c,d\geq 0$ and $a+b=c+d$, then $a^{k}+b^{k}=c^{k}+d^{k}$ forces $\{a,b\}=\{c,d\}$. For integer $k\geq 2$ we also prove a quantitative separation theorem on the hyperplane $a+b=c+d$, showing that distinct unordered pairs with the same sum $S$ produce values of $a^{k}+b^{k}$ that are separated by $\gg_{k}S^{k-2}$.
@@
Then any integer solution with shift $h$ must satisfy $M_{k}\mid h$; this modulus $M_{k}$ is the maximal squarefree modulus for which $x^{k}\equiv x\pmod{M_{k}}$ holds for all integers $x$. As immediate corollaries, if $k$ is prime then $k\mid h$, and for fixed $k$ the necessary condition $M_{k}\mid h$ leaves only a $1/M_{k}$ fraction of shifts. In particular, this creates a divisibility filter that eliminates, for instance, $99.96\%$ of all linear slices for $k=13$ (since $M_{13}=2730$). Combining MDO with our convexity analysis yields an exclusion zone principle. While this implies a combined bound on slice size $\min\{S,S+h\}\geq 2|h|/(k-1)$, we show that a global overlap argument provides a strictly stronger constraint. Finally, along any fixed slice $(S,h)$ we prove an asymptotic dominance bound $k\leq\max\{S,S+h\}\log 2$, beyond which no integer solutions can occur.
@@ 1 Introduction
For $k\geq 5$, however, the situation is dramatically different. As of December 2025, no non-trivial integer solutions to (1) are known in this range, and the existence of such solutions remains a difficult and largely open problem; see, for example, Browning [2] and the references therein, as well as Guy’s collection of unsolved problems [3].
@@
  • for integer $k\geq 2$, an exclusion zone of radius $\asymp\sqrt{\min\{S,S+h\}}$ around $\min\{S,S+h\}/2$ on the smaller-sum slice for each fixed $h\neq 0$;
@@
For integer solutions, combining MDO with the exclusion zone gives the lower bound $\min\{S,S+h\}\geq 2M_{k}/(k-1)$ whenever $h\neq 0$.
Organization of the paper. Section 2 develops the theory of the central slice $h=0$, including qualitative uniqueness and quantitative separation on $a+b=c+d$. Section 3 formulates central-slice uniqueness for general strictly convex functions. Section 4 proves the modular divisibility obstruction (MDO), its maximality, and corollaries (parity as a special case, density, prime exponents), and includes a table of $M_{k}$ for odd $k$. Section 5 contains the exclusion zone principle and its combination with MDO, yielding $\min\{S,S+h\}\geq 2M_{k}/(k-1)$ for $h\neq 0$, together with a separate “overlap” bound coming from the global ranges of $a^{k}+b^{k}$ on different slices. Section 6 establishes the asymptotic dominance bounds for fixed slices. Finally, Section 7 discusses heuristics and broader context.
Throughout, $\mathbb{N}=\{0,1,2,\dots\}$ and, unless explicitly stated otherwise, we take $a,b,c,d\in\mathbb{Z}_{\geq 0}$ and use $\log$ to denote the natural logarithm.
@@ 2.2 Quantitative separation on the central slice
Let $S\geq 2$ and put $n:=\lfloor S/2\rfloor$. For $t\in\{0,1,\dots,S\}$ set $F(t):=t^{k}+(S-t)^{k}$. Since F(t)=F(St)F(t)=F(S-t), every unordered pair $\{u,S-u\}$ is represented uniquely by $x:=\min\{u,S-u\}\in\{0,1,\dots,n\}$, and its value is $F(x)$. By Theorem 2.1, $F$ is strictly decreasing on $[0,S/2]$, hence the sequence $u_{t}:=F(t)$ for $t=0,1,\dots,n$ is strictly decreasing.
Apply Lemma 2.5 to $f_{S}$ on the arithmetic progression $0,1,\dots,n$. Then $\Delta_{t}:=u_{t+1}-u_{t}$ is strictly increasing in $t$, so $g_{t}:=u_{t}-u_{t+1}=-(\Delta_{t})$ is strictly decreasing in $t$. Therefore for any $0\leq x<y\leq n$ we have
$u_{x}-u_{y}=\sum_{t=x}^{y-1}g_{t}\ \geq\ g_{n-1}=u_{n-1}-u_{n}.$
Thus, for distinct unordered pairs with sum $S$,
$|a^{k}+b^{k}-c^{k}-d^{k}|\geq u_{n-1}-u_{n}.$
It remains to bound $u_{n-1}-u_{n}$ from below.
Case $S=2n$ (even). Then $u_{n}=2n^{k}$ and $u_{n-1}=(n-1)^{k}+(n+1)^{k}$, hence
$u_{n-1}-u_{n}$ $=(n+1)^{k}+(n-1)^{k}-2n^{k}$
$=2\sum_{\begin{subarray}{c}j\ \mathrm{even}\\ j\geq 2\end{subarray}}\binom{k}{j}n^{k-j}$
$\geq k(k-1)n^{k-2}.$
Case $S=2n+1$ (odd). Then $u_{n}=n^{k}+(n+1)^{k}$ and $u_{n-1}=(n-1)^{k}+(n+2)^{k}$, so
$u_{n-1}-u_{n}=(n+2)^{k}-(n+1)^{k}+(n-1)^{k}-n^{k}=\sum_{j=2}^{k}\binom{k}{j}n^{k-j}\bigl(2^{j}-1+(-1)^{j}\bigr).$
Here each coefficient $\bigl(2^{j}-1+(-1)^{j}\bigr)\geq 0$ and equals $4$ when $j=2$, hence
$u_{n-1}-u_{n}\geq 4\binom{k}{2}n^{k-2}=2k(k-1)n^{k-2}\geq k(k-1)n^{k-2}.$
Combining the cases yields
$|a^{k}+b^{k}-c^{k}-d^{k}|\geq k(k-1)\,n^{k-2}=k(k-1)\,\lfloor S/2\rfloor^{\,k-2}.$
Finally, for $S\geq 2$ one has $\lfloor S/2\rfloor\geq S/3$, hence $\lfloor S/2\rfloor^{k-2}\geq 3^{2-k}S^{k-2}$, giving the constant $C_{k}$. ∎
@@ 4 Modular constraints on the shift: the MDO
If $p\mid N$, then $x^{k}\equiv x\pmod{p}$ for all $x$. Restricting to units shows $x^{k-1}\equiv 1$ for all $x\in(\mathbb{Z}/p\mathbb{Z})^{\times}$, so the exponent of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ divides k1k-1. Since $(\mathbb{Z}/p\mathbb{Z})^{\times}$ is cyclic of order p1p-1, its exponent equals p1p-1. Hence $p-1\mid k-1$, i.e. $p\in\mathcal{P}_{k}$, and therefore $p\mid M_{k}$. As $N$ is squarefree, $N\mid M_{k}$. ∎
@@
For odd $k\in\{3,5,\dots,19\}$ we record $\mathcal{P}_{k}$, the squarefree modulus $M_{k}$, the density $1/M_{k}$, and the combined lower bound on the slice size (see Proposition 5.5) written as $2M_{k}/(k-1)$.
@@ 19 $\{2,3,7,19\}$ $798$ $1/798$ $1596/18$
While the table lists $M_{k}$ for $k\leq 19$, the value of $M_{k}$ grows very rapidly for exponents with highly composite k1k-1. For instance, at $k=61$ we have $60=2^{2}\cdot 3\cdot 5$, leading to
@@ 5 Geometry of shifted slices and exclusion zones
We will prove the following inequality, which shows that the $k$th power sum is bounded above by the $k$th power of its quadratic approximation at the mean $X$:
@@
Since the function $t\mapsto t^{k}$ is strictly increasing for $t\geq 0$, taking the $k$th root yields:
@@
For $m=0$, both sides are 1. For $m>\lfloor k/2\rfloor$, the LHS coefficient is 0, while RHS is non-negative.
Case $m=1$. LHS is $\binom{k}{2}=\frac{k(k-1)}{2}$. RHS is $\binom{k}{1}\frac{k-1}{2}=\frac{k(k-1)}{2}$. They are equal.
Case $m\geq 2$. We rewrite (4) (assuming $2m\leq k$):
@@
Since $m\geq 2$ and $0\leq j\leq m-1$, we have $k-m-j\leq k-2$ and $2m-j\geq m+1\geq 3$. Therefore
$\frac{k-m-j}{2m-j}\leq\frac{k-2}{3}\leq\frac{k-1}{2},$
so each factor in the product is $\leq(k-1)/2$. Hence
$\prod_{j=0}^{m-1}\frac{k-m-j}{2m-j}\leq\left(\frac{k-1}{2}\right)^{m},$
which proves (4).
@@
Proposition 5.5 (Combination with MDO: lower bound on slice size).
@@
If $h>0$, then $\delta\leq S/2$ and Theorem 5.4 gives $S^{2}/4\geq\delta^{2}\geq Sh/(2(k-1))$, hence $S\geq 2h/(k-1)$. If $h<0$, apply Theorem 5.4 to the pair $(c,d)$ with shift h>0-h>0: then $S+h\geq 2|h|/(k-1)$. Combining, $\min\{S,S+h\}\geq 2|h|/(k-1)$. The second bound follows since $M_{k}\mid h$ implies $|h|\geq M_{k}$ when $h\neq 0$. ∎
@@
It is important to note that the overlap bound (Lemma 5.2) is strictly stronger than the combined bound (Proposition 5.5) for all $k\geq 2$. That is,
@@ 6 Asymptotic dominance on fixed slices
Since $h\neq 0$, not all of $a,b,c,d$ lie in $\{0,1\}$; hence $M\geq 2$. If the maximal base $M$ appeared on both sides, then after reordering we may assume $a=c=M$. Cancelling $M^{k}$ gives $b^{k}=d^{k}$, hence $b=d$, and therefore h=(c+d)(a+b)=0h=(c+d)-(a+b)=0, a contradiction. Thus for $h\neq 0$ the value $M$ appears on exactly one side; say $M^{k}+b^{k}=c^{k}+d^{k}$ with $b,c,d\leq M-1$. Then $M^{k}-(M-1)^{k}\leq|d^{k}-b^{k}|\leq(M-1)^{k}$, which fails when $k>M\log 2$ by Lemma 6.1. Thus $k\leq M\log 2\leq S_{0}\log 2$. ∎
@@ 7 Concluding remarks
On the central slice $a+b=c+d$ we have complete uniqueness (Theorem 2.1) and quantitative separation (Theorem 2.6): distinct unordered pairs $\{a,b\}$ along $a+b=S$ produce values of $a^{k}+b^{k}$ separated by $\gg_{k}S^{k-2}$. This gives a simple local instance of sparsity heuristics for equal sums of two $k$th powers.
For shifted slices $h\neq 0$, MDO (Theorem 4.1) produces a squarefree modulus $M_{k}$ with $M_{k}\mid h$, maximal among such moduli by Lemma 4.2. This contains parity as a special case (Corollary 4.3), yields $k\mid h$ for prime $k$, and gives density $1/M_{k}$ of admissible shifts (Corollary 4.5). For many odd $k$ (e.g. $k=13$) this drastically thins the set of feasible $h$. Combined with the exclusion zone (Theorem 5.4) we obtain the lower bound $\min\{S,S+h\}\geq 2M_{k}/(k-1)$ (Proposition 5.5), which is an explicit structural constraint on the slice size. In addition, the overlap bound of Lemma 5.2 shows that the smaller of the two slice sums must satisfy $\min\{S,S+h\}\gg_{k}|h|$, giving a global geometric restriction based solely on the ranges of $a^{k}+b^{k}$ on the two slices. Finally, along a fixed slice $(S,h)$, the dominance bound $k\leq\max\{S,S+h\}\log 2$ (Theorem 6.2) shows that only finitely many exponents can occur.